OpenStudy (anonymous):

#1. Determine if the relation is a function. Explain your reasoning. [(4,5),(3,-2),(-2,5),(4,7)] #2 The function h(t) = -16t2 + 68t + 2 represents the height h(t) of a football in feet t seconds after it is kicked. a. Explain how to use the graph to find h(2). b. Find the value of h(2) algebraically. c. Determine if the value of h(2) is in the domain given the situation. Why or why not. d. Find the value of h(3.5) algebraically. e. Determine if the value of h(3.5) is in the domain given the situation. Why or why not. f. Find the value of h(4.5) algebraically. g. Dete

OpenStudy (mertsj):

A function is a set of ordered pairs in which no two ordered pairs have the same first number.

OpenStudy (mertsj):

a. Find the y value that corresponds to an x value of 2. b. h(2)=-16(2)^2+68(2)+2=7 c. Yes 2 is in the domain. It gives a real value for the height of the ball after 2 seconds. d. h(3.5)=-16(3.5)^2+68(3.5)+2=44 e. It is in the domain. It gives a real value for the height of the ball. f. h(4.5)= -16(4.5)^2+68(4.5)+2=-16

OpenStudy (anonymous):

g. Determine if the value of h(4.5) is in the domain given the situation. Why or why not. h. Define the domain given the situation. i. Define the range give the situation. #3. Given the equation -20x + ¼ y = 10 a. Find the slope. b. Find the y-intercept #4. Find the value of x so that the line (x,6) and (10,-3) has a slope of -3/2. Then write the equation of the line in y-intercept form. #5. On a map, Kester Drive passes through R(4,-11) and S(0,-9) and Randolph Road passes through J(6,-2) and K(4,-5). If they are straight lines, are the two streets perpendicular? Explain.

OpenStudy (anonymous):

Of coarse f(4.5) is in the domain given it can be graphed.

OpenStudy (anonymous):

The domain is the set of real numbers that can be graphed given the situation. This includes nothing but positive integers.

OpenStudy (anonymous):

#3. The slope is always next to x. Therefore m = -20. The y-intercept is the number that is solely by itself, which is 10. #4. To find the slope, we subtract the y's and then we subtract the x's and divide the two. We take x -10 and 6--3. This gives us x -10 and 9. We divide both numbers. 9/x-10 is equal to -3/2.